OpenStudy (anonymous):

Find the angle between the given vectors to the nearest tenth of a degree. u = <2, -4>, v = <3, -8>

OpenStudy (zzr0ck3r):

\[a\cdot b=|a|*|b|*cos(\theta), where\space \theta\space is\space the\space angle\space between\space the\space vectors\]

OpenStudy (anonymous):

I tried everything, but I'm not getting the answer I need...

OpenStudy (anonymous):

\[ \cos \theta = \frac {a \cdot b} {|a| |b| } \] take the arccos to find \[ \theta\]

OpenStudy (anonymous):

I know that's how you solve it, however my issue is that I am not getting the right answer. Obviously, I'm doing something wrong.

OpenStudy (zzr0ck3r):

\[a \cdot b = 38\\ |a|*|b| = \sqrt{4+16}*\sqrt{73}=2\sqrt{5}\sqrt{73 }\\\theta=\cos^{-1}(\frac{19}{\sqrt{5}\sqrt{73}})\approx0.10487\]

OpenStudy (zzr0ck3r):

@Loveiskey18 do you understand?

OpenStudy (anonymous):

\[ \Large \cos \theta = \frac {a \cdot b} {|a| |b| }\\ \Large \cos \theta = \frac {(2*3)+(-4*-8)} {\sqrt{2^2+3^2}*\sqrt{4^4+8^2} }\\ \]

OpenStudy (zzr0ck3r):

your vectors are wrong on the bottom @cinar shuold be \[\sqrt{2^2+4^2}\\and\\\sqrt{3^2+8^2}\]

OpenStudy (anonymous):

@zzr0ck3r Sorry I took so long, I was attempting in understanding this but I am not. I just keep getting stuck at the part where you input the vectors.

OpenStudy (anonymous):

you are right

OpenStudy (zzr0ck3r):

where I input the vectors? do you know what the dot product is, and the TWO different ways it is defined?

OpenStudy (zzr0ck3r):

you need to know this before you can answer this.

OpenStudy (zzr0ck3r):

@Loveiskey18 this will go allot faster if you answer:)

OpenStudy (anonymous):

I'm sorry, I was attempting it on my own...again

OpenStudy (zzr0ck3r):

ok just stay with me and ill walk you through it?

OpenStudy (anonymous):

My ending results: arccos (38/√(13)√(-80))

OpenStudy (zzr0ck3r):

nope

OpenStudy (zzr0ck3r):

can you tell me what \[<2,-4>\cdot<3,-8>=?\]

OpenStudy (anonymous):

arccos (38/√(-1040)

OpenStudy (zzr0ck3r):

square root of negative number is the first thing that I can see is wrong

OpenStudy (zzr0ck3r):

please just stick with me and stop posting answers:)

OpenStudy (anonymous):

Laughing out loud, okay.

OpenStudy (zzr0ck3r):

\[<2,-4>\cdot<3,-8>=?\]

OpenStudy (zzr0ck3r):

I think you did this part already, its 38 right?

OpenStudy (anonymous):

Yes.

OpenStudy (zzr0ck3r):

do you understand \[\theta=cos^{-1}(\frac{a\cdot b}{|a||b|})\]

OpenStudy (anonymous):

Yes.

OpenStudy (zzr0ck3r):

so we need |a|*|b| correct?

OpenStudy (anonymous):

Correct.

OpenStudy (zzr0ck3r):

\[a=<2,-4> \\|a| =\sqrt{2^2+(-4)^2}=\sqrt{4+16}=\sqrt{20}=\sqrt{4*5}=\sqrt{4}\sqrt{5}=2\sqrt{5}\\correct?\]

OpenStudy (anonymous):

Yes.

OpenStudy (zzr0ck3r):

\[b=<3,-8>\\|b|=\sqrt{3^2+(-8)^2}=\sqrt{9+64}=\sqrt{73}\]correct?

OpenStudy (anonymous):

Yes.

OpenStudy (zzr0ck3r):

if there is anything you dont understand just tell me, I am in no rush

OpenStudy (anonymous):

I acknowledge everything completely.

OpenStudy (zzr0ck3r):

ok then we have\[\theta=\cos^{-1}(\frac{a\cdot b}{|a|*|b|})=cos^{-1}(\frac{38}{2\sqrt{5}\sqrt{73}})=\cos^{-1}(\frac{19}{\sqrt{5*73}})=\cos^{-1}(\frac{19}{\sqrt{365}})\]

OpenStudy (zzr0ck3r):

\[\approx0.10487\]

OpenStudy (anonymous):

To be honest I do understand all of this, and actually I already obtained that answer before I asked for help, however I wasn't sure. The problem is... that is not the answer I am looking for, well it is. I just do not know how to convert it into degrees or the appropriate answer: My choices: 3.0° 6.0° -7.0° 16.0° Nonetheless, I am very much grateful for you putting in the time and energy to help me.

OpenStudy (zzr0ck3r):

we got radians as the answer\[radian\space answer*\frac{180}{\pi}=degree\space answer\]

OpenStudy (zzr0ck3r):

\[.10487*\frac{180}{\pi}\approx6.0081\approx6\]

OpenStudy (zzr0ck3r):

next time you are confused about radians to degree imagine if you have pi, then you know that is 180 degrees well what can we multiply pi by to get 180? pi(180/pi) = 180

OpenStudy (zzr0ck3r):

and if you have 180 degrees you know that is pi well what can we multiply 180 by to get pi? 180*(pi/180)=pi

OpenStudy (zzr0ck3r):

ok so, after all that....are we good to go now?

OpenStudy (anonymous):

Thank You so much :D Do you mind verifying a couple questions for me, please. Just tell me if they are right or wrong.